# Tool-LMM: A Large Multi-Modal Model for Tool Agent Learning

### v1

@@ Abstract
Recently, the astonishing performance of large language models (LLMs) in natural language comprehension and generation tasks triggered lots of exploration of using them as central controllers to build agent systems. Multiple studies focus on bridging the LLMs to external tools to extend the application scenarios. However, the current LLMs’ perceiving tool-use ability is limited to a single text query, which may result in ambiguity in understanding the users’ real intentions. LLMs are expected to eliminate that by perceiving the visual- or auditory-grounded instructions’ information. Therefore, in this paper, we propose Tool-LMM, a system incorporating open-source LLMs and multi-modal encoders so that the learnt LLMs can be conscious of multi-modal input instruction and then select the function-matched tool correctly. To facilitate the evaluation of the model’s capability, we collect a dataset featured by consisting of multi-modal input tools from HuggingFace. Another important feature of our dataset is that our dataset also contains multiple potential choices for the same instruction due to the existence of identical functions and synonymous functions, which provides more potential solutions for the same query. The experiments reveal that our LMM is capable of recommending appropriate tools for multi-modal instructions. Codes and data are available at https://github.com/Tool-LMM/Tool-LMM.
@@ 1 Introduction
Figure 1: Comparisons of results outputted by tool agent with different inputs. The individual posed the same question thrice. Initially, no image was provided, causing the confusion. In the subsequent instances, images from varied scenes were uploaded, and our Tool-LMM can correctly choose the correct tool based on the multi-modal inputs and generate distinct responses based on the differing pictorial contexts.
@@
To remedy the previous works only accepting single text instructions, in this paper, we introduce a novel system, Tool-LMM, integrating multi-modal encoders with open-source LLMs to synthesize multi-modal information for correct external tool identification. The distinction between our work and previous works is delineated in Table 1. Compared to previous open-sourced works, our system is the first attempt to train a large multi-modal model for tool agent learning. In light of the fact that existing tool learning benchmarks  [15, 12, 39, 23, 21, 37, 42, 27], do not explicitly construct text-to-any-modality pairs for alleviating text ambiguity, which cannot facilitate the training of our Tool-LMM, in this paper, we collect a novel dataset, named ToolMMBench from the HuggingFace platform, one of the largest machine learning communities containing thousands of APIs. It is also worth noting that our ToolMMBench supports multi-options for each query, as demonstrated in Table 1. This feature is more closely related to real-life scenarios, where multiple alternatives exist to address the same challenges, albeit with varying performance levels. Such a feature would also make our dataset suitable for other applications in future, such as the study of integrating multiple outputs from different tools to better feedback to the user.
@@
  • We develop Tool-LMM, a multi-modal tool agent system that can perceive the visual- or audio-grounded instructions’ information via incorporating multi-modal encoders and open-source large language models.
@@
  • We design some evaluation metrics and conduct experiments with extensive subset ablations on multiple popular LLMs, and by fine-tuning Tool-LMM on ToolMMBench, we observe a promising accuracy, 88.19% of the tool selection. which demonstrates the effectiveness of our method.
@@ 2 Related Works
Multi-modal Large Language Models. LLMs have increasingly become the cornerstone of many natural language processing (NLP) applications. Subsequently, many efforts, including GPT-4V  [19], LLaVA  [16], BLIP-2  [14], Minigpt-4  [46] developed the LLM to vision space by aligning off-the-shelf frozen pre-trained image encoders. And similar ideas are generalized to video and audio modalities, like NeXTGPT  [38], and VideoLLaMA  [44]. Our proposed system, Tool-LMM, focuses on exploring tool usage awareness and selection. Our goal is to empower the LLM’s ability to be conscious of the external tools’ function and select the suitable tools, to act as a controller, instead of enhancing the ability of multi-modal understanding and reasoning.
@@ 3.1 API Collection
For each model, we document the model card and extract the key information, including the domain, api_\_name, api_\_call, parameters, coarse_\_functionality, fine_\_functionality, example codes, and descriptions. Further, to prevent an overabundance of APIs with identical functionalities leading to excessive options for subsequent instruction matching (e.g., in the imaging realm, where two APIs might differ only in their network architectures but are identical in terms of the datasets they were trained on and their functionalities), we adopt a strategy based on the API’s download statistics. For any single functionality category, we retain a maximum of five of the most downloaded APIs.
@@ 3.4 Dataset Statistics
4 Tool-LMM
We introduce Tool-LMM, a method that leverages the LLM performing the role of multi-modality tool agent. The architecture of this network is depicted in Figure 4. Considering the diversity of our data, which includes images, videos, and audios, we incorporate ImageBind  [9] as our primary multimodal encoder. ImageBind  [9] is designed to create a unified embedding space that integrates six distinct modalities, utilizing the advancements in large-scale vision-language models.
For the large language model component of Tool-LMM, we select a range of leading LLMs, including Vicuna  [5], Llama  [33], Llama2  [34], and Llama2-Chat  [34]. These models are chosen for their exemplary performance in various areas, such as inference capabilities, knowledge retention, and comprehension skills. Consequently, our system is equipped with the proficiency to interpret multi-modal input instructions effectively.
@@ 5 Evaluation Metrics
The testing set is divided into several subsets based on the three attributes. For a subset $\mathbf{S}$, supposing containing $n$ instructions $\{x_{1},x_{2},x_{3}\dots x_{n}\}$, where each instruction $x_{i}$ matches an API list $y_{i}$, which consists of a list of $M$ suitable APIs $A_{i}^{1},A_{i}^{2},A_{i}^{3},\dots,A_{i}^{M}]$. Thus, the collection of API lists corresponding to this instruction subset combined to a set $\mathbf{Y}$, $\{y_{1},y_{2},y_{3}\dots y_{n}\}$. With these symbols defined, assuming the subset $\mathbf{S}$ yields an API set $\{z_{1},z_{2},z_{3}\dots z_{n}\}$, output from our model, we can represent the accuracy of each mentioned subset using the following formula:
@@ 6 Experiments
In this section, we conduct experiments on our collected dataset, evaluating the performance of our system Tool-LMM. We start by introducing the experiment setup in Sec. 6.1 and data processing in Sec. 6.2. In Sec. 6.3, we assess Tool-LMM’s ability on our benchmark, followed by results analysis in Sec. 6.4. And we conduct qualitative experiment and visualize the result in Sec. 6.5.
@@ 6.2 Data processing
Table 5: Experimental results of ToolLMM. We present the performance of eight large language models with different configurations, and we report the performance at different epochs of training to reflect the changing trend.
@@ 6.3 Experimental Results
Overall, the results affirm that our system Tool-LMM effectively brings out the tool-use potential in LLMs and enhances the LLMs to deal with multi-modal instruction inputs.
@@ 6.4 Results Analysis
Figure 5: Qualitative example of Tool_LMM on text ambiguity case. Tool_LMM firstly receives the text instruction and perceives information from the input image to determine the corresponding task, select the appropriate API that meets the requirements from the API corpus, and then call the API to execute, and finally get the quality-improved image. Zoom in for the best view.
@@ 6.5 Visualization
We visualize qualitative result of Tool_LMM, to verify its effectiveness. As shown in Figure 5, our system is asked to improve the quality of the input image without expaining the reason for the poor image quality, and Tool_LMM is capable of perceiving the multi-modal information and choosing the correct API by combining the instruction and visual information, which proves that our Tool_LMM can effectively accomplish the users’ needs.
@@ 7 Conclusion
In this work, we introduce how to empower the tool-augmented LLMs with perceiving visual- or audio-grounded instructions’ information. We collect a benchmark, ToolMMBench, from the HuggingFace Platform, which covers 932 high-quality machine learning related APIs. The finetuning of this dataset on various popular LLMs has yielded promising results, demonstrating that Tool-LMM can accurately recommend appropriate tools for multi-modal instructions.
@@ 8.4 Prompt Construction
### User: You are an NLP task expert. Given an API, you need to generate 20 different user queries that utilize the API function, adhering to the following input and output format for each query: Input: API Name: This is the name of the API Function. Description: This is a detailed description of the model. Prohibit Words: These are words that are prohibited from appearing in the output. Output: Query1: This is an instruction that can utilize the API function. Query2: This is an instruction that can utilize the API function. Query3: This is an instruction that can utilize the API function. Query4: This is an instruction that can utilize the API function. Query5: This is an instruction that can utilize the API function. Query20: This is an instruction that can utilize the API function. Below are some examples: Example 1: Input: API Name: microsoft/codereviewer Description: CodeReviewer is a model pre-trained with code change and code review data to support code review tasks. Prohibit Words: ”API, tools, model” Output: Query1: I’ve made some changes to my Python code. Can you review this snippet for me? ”def addNums(a, b): return a + b” Query2: I’m not very confident in my Java coding skills. Could you check this piece for potential issues? ”public int divide(int x, int y) { return x / y; }” Query3: I’m trying to refactor this C# method. Does it look okay? ”public void PrintName(string name) { Console.WriteLine(name); }” Query4: I wrote this SQL query, and I’m unsure if it’s optimized. Can you review it? ”SELECT * FROM users WHERE age $\geq$ 25;” Query5: Here’s a piece of JavaScript function I came up with. Any suggestions for improvement? ”function greet(name) { return ’Hello ’ + name; }” Query6: I’ve been learning Ruby recently. Would appreciate your thoughts on this code segment: ”def multiply(x, y) x * y end” Query7: I’m not sure if this CSS is correct. Can you review it? ”h1 { color: blue; font-size: 30px; }” Query8: I just started with Swift. Here’s a function I wrote. Could you review it for best practices? ”func displayMessage(message: String) { print(message) }” Query9: My colleague wrote this PHP function, but it seems off. Can you review it for me? ”function subtract(x,x,y) { return $x$-y; }” Query20: I’m learning Kotlin and wrote this simple method. Mind taking a look? ”fun sum(a: Int, b: Int): Int = a + b” Example 2: Input: API Name: Salesforce/blip-vqa-base To be continued
Continued
Description: This model is a base-sized ViT-based BLIP trained on the visual question answering task VQA2.0 dataset. The model’s input is an image and a text question, and the model’s output is a text answer.
@@
Prohibit Words: ”API, tools, model, VQA2.0” Output: Query1: I recently visited a new city and took some photographs. Based on this particular image, can you tell me what is the weather conditions like? Query2: I was at a photo exhibition and took this picture. Could you provide information about was this taken in a Latin American country judging from the dress? Query3: I was at the park and saw a dog playing. By analyzing this photo, can you tell if the frisbee it is using is hard or soft? Query4: My cousin went to a new food joint and ordered their specialty hot dog. By looking at this picture, can you determine if the hot dog is larger than a normal one? Query5: I attended a family gathering and took a group photo. By analyzing the photograph, how many kids can you see in the picture? Query6: Looking at this captivating portrait, I noticed an object in the individual’s hand. It seems to be a sweet treat. What is this person holding? Query7: I was at the beach and took a photo of a man skiing. By examining the image, can you tell if the wave is chasing him? Query8: I visited the local train station and captured an image of a passing train. Can you identify what color the train is from the picture? Query9: This cityscape presents an old-fashioned vehicle that stands out amidst the modern structures. I’d like to know more about this vehicle. Could you tell me its model and color?” Query20: Given the serene backdrop of sailboats and a calm sea, can you tell me how many boats are there? Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
@@
Continued
Query5: Craft a depiction of a magician performing tricks in front of an amazed audience from my image.
Query6: Could you reshape this photograph into a bustling train station with intricate ironwork and arches? Query7: Transform this snapshot into a charming cottage nestled among rolling hills and wildflowers. Query8: I’m curious about how this picture would be reimagined as a bustling harbor with ships, cranes, and cargo containers. Query9: Use my image to depict a family picnic in a sunlit park with children playing. Query20: I’d love to see this image as a gardener carefully pruning roses in a blooming garden. Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response; 4. The queries should not convey or imply multimodal information. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]

### v2

# MLLM-Tool: A Multimodal Large Language Model For Tool Agent Learning

@@ Abstract
Recently, the astonishing performance of large language models (LLMs) in natural language comprehension and generation tasks triggered lots of exploration of using them as central controllers to build agent systems. Multiple studies focus on bridging the LLMs to external tools to extend the application scenarios. However, the current LLMs’ perceiving tool-use ability is limited to a single text query, which may result in ambiguity in understanding the users’ real intentions. LLMs are expected to eliminate that by perceiving the visual- or auditory-grounded instructions’ information. Therefore, in this paper, we propose MLLM-Tool, a system incorporating open-source LLMs and multi-modal encoders so that the learnt LLMs can be conscious of multi-modal input instruction and then select the function-matched tool correctly. To facilitate the evaluation of the model’s capability, we collect a dataset featured by consisting of multi-modal input tools from HuggingFace. Another important feature of our dataset is that our dataset also contains multiple potential choices for the same instruction due to the existence of identical functions and synonymous functions, which provides more potential solutions for the same query. The experiments reveal that our MLLM-Tool is capable of recommending appropriate tools for multi-modal instructions. Codes and data are available at https://github.com/MLLM-Tool/MLLM-Tool.
@@ 1 Introduction
Figure 1: Comparisons of results outputted by tool agent with different inputs. The individual posed the same question thrice. Initially, no image was provided, causing the confusion. In the subsequent instances, images from varied scenes were uploaded, and our MLLM-Tool can correctly choose the correct tool based on the multi-modal inputs and generate distinct responses based on the differing pictorial contexts.
@@
To remedy the previous works only accepting single text instructions, in this paper, we introduce a novel system, MLLM-Tool, integrating multi-modal encoders with open-source LLMs to synthesize multi-modal information for correct external tool identification. The distinction between our work and previous works is delineated in Table 1. Compared to previous open-sourced works, our system is the first attempt to train a large multi-modal model for tool agent learning. In light of the fact that existing tool learning benchmarks  [15, 12, 39, 23, 21, 37, 42, 27], do not explicitly construct text-to-any-modality pairs for alleviating text ambiguity, which cannot facilitate the training of our MLLM-Tool, in this paper, we collect a novel dataset, named ToolMMBench from the HuggingFace platform, one of the largest machine learning communities containing thousands of APIs. It is also worth noting that our ToolMMBench supports multi-options for each query, as demonstrated in Table 1. This feature is more closely related to real-life scenarios, where multiple alternatives exist to address the same challenges, albeit with varying performance levels. Such a feature would also make our dataset suitable for other applications in future, such as the study of integrating multiple outputs from different tools to better feedback to the user.
@@
  • We develop MLLM-Tool, a multi-modal tool agent system that can perceive the visual- or audio-grounded instructions’ information via incorporating multi-modal encoders and open-source large language models.
@@
  • We design some evaluation metrics and conduct experiments with extensive subset ablations on multiple popular LLMs, and by fine-tuning MLLM-Tool on ToolMMBench, we observe a promising accuracy, 88.19% of the tool selection. which demonstrates the effectiveness of our method.
@@ 2 Related Works
Multi-modal Large Language Models. LLMs have increasingly become the cornerstone of many natural language processing (NLP) applications. Subsequently, many efforts, including GPT-4V  [19], LLaVA  [16], BLIP-2  [14], Minigpt-4  [46] developed the LLM to vision space by aligning off-the-shelf frozen pre-trained image encoders. And similar ideas are generalized to video and audio modalities, like NeXTGPT  [38], and VideoLLaMA  [44]. Our proposed system, MLLM-Tool, focuses on exploring tool usage awareness and selection. Our goal is to empower the LLM’s ability to be conscious of the external tools’ function and select the suitable tools, to act as a controller, instead of enhancing the ability of multi-modal understanding and reasoning.
@@ 3.1 API Collection
For each model, we document the model card and extract the key information, including the domain, api__\__name, api__\__call, parameters, coarse__\__functionality, fine__\__functionality, example codes, and descriptions. Further, to prevent an overabundance of APIs with identical functionalities leading to excessive options for subsequent instruction matching (e.g., in the imaging realm, where two APIs might differ only in their network architectures but are identical in terms of the datasets they were trained on and their functionalities), we adopt a strategy based on the API’s download statistics. For any single functionality category, we retain a maximum of five of the most downloaded APIs.
@@ 3.4 Dataset Statistics
4 MLLM-Tool
We introduce MLLM-Tool, a method that leverages the LLM performing the role of multi-modality tool agent. The architecture of this network is depicted in Figure 4. Considering the diversity of our data, which includes images, videos, and audios, we incorporate ImageBind  [9] as our primary multimodal encoder. ImageBind  [9] is designed to create a unified embedding space that integrates six distinct modalities, utilizing the advancements in large-scale vision-language models.
For the large language model component of MLLM-Tool, we select a range of leading LLMs, including Vicuna  [5], Llama  [33], Llama2  [34], and Llama2-Chat  [34]. These models are chosen for their exemplary performance in various areas, such as inference capabilities, knowledge retention, and comprehension skills. Consequently, our system is equipped with the proficiency to interpret multi-modal input instructions effectively.
@@ 5 Evaluation Metrics
The testing set is divided into several subsets based on the three attributes. For a subset $\mathbf{S}$, supposing containing $n$ instructions $\{x_{1},x_{2},x_{3}\dots x_{n}\}$, where each instruction $x_{i}$ matches an API list $y_{i}$, which consists of a list of $M$ suitable APIs $[A_{i}^{1},A_{i}^{2},A_{i}^{3},\dots,A_{i}^{M}]$. Thus, the collection of API lists corresponding to this instruction subset combined to a set $\mathbf{Y}$, $\{y_{1},y_{2},y_{3}\dots y_{n}\}$. With these symbols defined, assuming the subset $\mathbf{S}$ yields an API set $\{z_{1},z_{2},z_{3}\dots z_{n}\}$, output from our model, we can represent the accuracy of each mentioned subset using the following formula:
@@ 6 Experiments
In this section, we conduct experiments on our collected dataset, evaluating the performance of our system MLLM-Tool. We start by introducing the experiment setup in Sec. 6.1 and data processing in Sec. 6.2. In Sec. 6.3, we assess MLLM-Tool’s ability on our benchmark, followed by results analysis in Sec. 6.4. And we conduct qualitative experiment and visualize the result in Sec. 6.5.
@@ 6.2 Data processing
Table 5: Experimental results of MLLM-Tool. We present the performance of eight large language models with different configurations, and we report the performance at different epochs of training to reflect the changing trend.
@@ 6.3 Experimental Results
Overall, the results affirm that our system MLLM-Tool effectively brings out the tool-use potential in LLMs and enhances the LLMs to deal with multi-modal instruction inputs.
@@ 6.4 Results Analysis
Figure 5: Qualitative example of MLLM-Tool on text ambiguity case. MLLM-Tool firstly receives the text instruction and perceives information from the input image to determine the corresponding task, select the appropriate API that meets the requirements from the API corpus, and then call the API to execute, and finally get the quality-improved image. Zoom in for the best view.
@@ 6.5 Visualization
We visualize qualitative result of MLLM-Tool, to verify its effectiveness. As shown in Figure 5, our system is asked to improve the quality of the input image without expaining the reason for the poor image quality, and MLLM-Tool is capable of perceiving the multi-modal information and choosing the correct API by combining the instruction and visual information, which proves that our MLLM-Tool can effectively accomplish the users’ needs.
@@ 7 Conclusion
In this work, we introduce how to empower the tool-augmented LLMs with perceiving visual- or audio-grounded instructions’ information. We collect a benchmark, ToolMMBench, from the HuggingFace Platform, which covers 932 high-quality machine learning related APIs. The finetuning of this dataset on various popular LLMs has yielded promising results, demonstrating that MLLM-Tool can accurately recommend appropriate tools for multi-modal instructions.
@@ 8.4 Prompt Construction
[…]
@@
Continued Description: This model is a base-sized ViT-based BLIP trained on the visual question answering task VQA2.0 dataset. The model’s input is an image and a text question, and the model’s output is a text answer. Prohibit Words: ”API, tools, model, VQA2.0” Output: Query1: I recently visited a new city and took some photographs. Based on this particular image, can you tell me what is the weather conditions like? Query2: I was at a photo exhibition and took this picture. Could you provide information about was this taken in a Latin American country judging from the dress? Query3: I was at the park and saw a dog playing. By analyzing this photo, can you tell if the frisbee it is using is hard or soft? Query4: My cousin went to a new food joint and ordered their specialty hot dog. By looking at this picture, can you determine if the hot dog is larger than a normal one? Query5: I attended a family gathering and took a group photo. By analyzing the photograph, how many kids can you see in the picture? Query6: Looking at this captivating portrait, I noticed an object in the individual’s hand. It seems to be a sweet treat. What is this person holding? Query7: I was at the beach and took a photo of a man skiing. By examining the image, can you tell if the wave is chasing him? Query8: I visited the local train station and captured an image of a passing train. Can you identify what color the train is from the picture? Query9: This cityscape presents an old-fashioned vehicle that stands out amidst the modern structures. I’d like to know more about this vehicle. Could you tell me its model and color?” Query20: Given the serene backdrop of sailboats and a calm sea, can you tell me how many boats are there? Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
@@
Continued Query5: Craft a depiction of a magician performing tricks in front of an amazed audience from my image. Query6: Could you reshape this photograph into a bustling train station with intricate ironwork and arches? Query7: Transform this snapshot into a charming cottage nestled among rolling hills and wildflowers. Query8: I’m curious about how this picture would be reimagined as a bustling harbor with ships, cranes, and cargo containers. Query9: Use my image to depict a family picnic in a sunlit park with children playing. Query20: I’d love to see this image as a gardener carefully pruning roses in a blooming garden. Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response; 4. The queries should not convey or imply multimodal information. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]